OpenStudy (anonymous):

A polyhedron has 7 faces and 12 edges. What is the number of vertices of the polyhedron? Answer 3 5 7 9

Directrix (directrix):

Vertices + Faces = Edges + 2 V + 7 = 12 + 2 V = 7

OpenStudy (anonymous):

7

OpenStudy (anonymous):

thanks